OpenStudy (anonymous):

Find a number whose product with 8 is the same as its sum with 49. (Hint: you may want to try guess and check or use a table.) pls helppp

OpenStudy (kirbykirby):

If I understand the problem correctly, they want a number (let's call it "x") such that multiplying x with 8 will give the same as adding x with 49?

OpenStudy (kirbykirby):

So we want "8x" (8 times x) to be the same thing as "49 + x" So 8x = 49 + x 8x - x = 49 7x = 49 x = 7, so 7 is the number